OpenStudy (muzzack):

An equation is shown below: 1 over 2 multiplied by x plus 3 over 2 equals 2 to the power of x What is the solution to the equation? x = 1 x = 2 x = 4 x = 8

OpenStudy (muzzack):

\[\frac{ 1 }{ 2 }x + \frac{ 3 }{ 2 } = 2^{x}\]

OpenStudy (muzzack):

@AccessDenied @anonymous_user @alias @AydenCarpenter14 @anqie37 @AmorVida @mathmale @TheRealMeeeee @TSwizzle @taffalousswagg_bru @harrypotter2 @harrypotter2 @ulaa. @inkyvoyd @KendrickLamar2014 @ltp1212 @msumner @hippo4 @bbbbbbbbbbbbbbbbbb @Barrelracer011

OpenStudy (muzzack):

@Hero

hero (hero):

Try plugging in each value one at a time until both sides agree.

hero (hero):

Start with x = 1

OpenStudy (anonymous):

well, you can try "simplifying" it by multiplying two one both sides You'll get x + 3 = 2^(x+1)

OpenStudy (muzzack):

well it gave me for 2 \[x = \frac{ 5 }{ 2 }\]

hero (hero):

Did you try x = 1 yet?

OpenStudy (muzzack):

2 =2

OpenStudy (muzzack):

thats infinite solution

hero (hero):

2 = 2 means that both sides agree and that x = 1 is the correct solution.

OpenStudy (muzzack):

oh ok thx @Hero you saved the day hehe lol